# Twin electroweak bubble nucleation and gravitational wave under the $S_{3}$ symmetry of two-Higgs-doublet model

### v4

@@ I INTRODUCTION
Particle physics today not only studies each individual problem in depth but also combines it with Cosmology and Astronomy. The baryon asymmetry is a problem of intersection among those fields. The possible causes of this problem are summarized and presented quite extensively. Indeed, for Higgs 125 GeV, Standard Model (SM) has only a second order electroweak phase transition mkn ; SME ; SMEb ; SMEc ; SMEd ; michela . Therefore, SM should be extended (for example, Refs. plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; chr ; cde ; kusenko ). There are 3 main branches: the first is due to the existence of exotic particles (beyond the standard model (SM)); the second is because of parameters or interactions in SM changing at a electroweak transition temperature plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; chr ; cde ; kusenko ; the third be having more than 4 dimensions of spacetime pn .
@@
  • Furthermore, in Ref. dori ; 124 ; 127 there are many scenarios among the additional Higgs bosons were studied in the EWPT problem, specifically $m_{H^{\pm}}=m_{A}$ or $m_{H}=m_{H^{\pm}}$.
@@ II.1 The Higgs potential
$\begin{split}&m_{H^{\pm}}^{2}=-l_{2}v^{2},\\ &m_{A}^{2}=-\dfrac{1}{2}(2l_{2}-l_{3})v^{2},\\ &m_{H}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}+\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{H}}{4}.v^{2},\\ &m_{h}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}-\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{h}}{4}.v^{2}\,.\end{split}$ (4)
@@ II.2 The temperature effective Higgs potential and cotronling EWPT
With $v_{1}$ different from $v_{2}$, this model has two stages of phase transition and assume $v_{1}<v_{2}$, which means $21<a<$ 2 or $1<\tan\beta$. As the above sections pointed out, in 2HDM-$S_{3}$, the masses of particles can be changed such that there are no mixing terms between $v_{1}$ and $v_{2}$. So the correct effective potential for this model is
$V_{eff}^{S_{3}}=V_{eff}^{S_{3}}(v_{1})+V_{eff}^{S_{3}}(v_{2}).$ (5)
@@
From many conclusions, we propose a way to investigate or control the Higgs potential to have a first order phase transition or $S=\frac{v_{c}}{T_{c}}>1$ ($v_{c}$ and $T_{c}$ are VEV and critical temperature respectively),
  • Build high temperature effective potential. Select the investigation variable, usually the masses of new particles.
  • Survey in the absence of daisy loops. Specify the range of values for variables to have $S>1$.
  • If the masses of new particles are greater than 200 GeV, there are no need to use additional daisy loops.
@@
According to Fig.2 and 3, the two potential functions corresponding to $v_{2}$ and $v_{1}$ get closer to each other as $a$ increases. With the uniform convergence of second minima (i.e. always less than $246$ GeV), the two effective potentials are always similar to each other. Although Fig.3 only corresponds to one mass pair, it is quite general, other mass pairs are presented in Appendix A.
Figure 2: The effective potential $V_{eff}$ at $T=T_{c}$. $m_{H^{\pm}}(v_{2})=m_{A}(v_{2})=500$ GeV, $m_{H}(v_{2})=250$ GeV. The red line corresponds to $V_{eff}(v_{2})$. The orange line corresponds to $V_{eff}(v_{1})$.
@@
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-14}$
@@ III.2 Gravitational waves
The next, by using a possible approximation phong2022 , because when calculating GW we only consider the effective potential and scalar field component 71 ; phu5gw ,
$\frac{\beta}{H^{*}}\approx\left[T\frac{d\left(0.45\frac{E^{T}_{sph}}{T}\right)}{dT}\right]_{T=T_{C}},$ (24)
@@
The gravitational wave detection frequency is solved in the range below a few mHz. At such low frequencies there is currently no experimental data. However, future detectors may be able to detect gravitational energy densities in the small frequency range as summarized in Table 4. The LISA and DECIGO detectors can pick up the gravitational wave signal generated by the EWPT process in the 2HDM-$S_{3}$ model as shown in Fig. 6.
@@ IV CONCLUSION AND OUTLOOKS
According to the data tables, at the same value of $a$, the ratio $E_{b}/v_{b}$ is always equal. By the linear scaling law of sphaleron energy scaling1 ; scaling2 ; scaling3 , $E_{b}/v_{b}=E_{b0}/v_{b0}=const$ ($E_{b0}$ is the sphaleron energy at $T=0$, Eq. (13) calculated by replacing the effective potential at temperature $T$ with one at 0), the evolution (increase or decrease of sphaleron energy) in the two phase transitions is the same. However, $a$ and this progression increase proportionally.
Additionally to test the scaling law (however, this was initially only an approximation to solve the Sphaleron energy problem scaling2 ; scaling3 ). We estimate some cases like Table 5, $E_{b}/v_{b}$ and $E_{b0}/v_{b0}$ differ by no more than $24\%$. Thus the results are not very compatible with the scaling law or this law is only true in a non-zero temperature range but is not true in the temperature range from 0 to $T_{c}$. For some other cases the calculation also breaks this law as in Ref.phascaling . Because there are many triggers outside SM and energy scales other than SM. Furthermore, the solutions $h(\xi)$ and $f(\xi)$ in Eqs. (13) will approach $1$ very slowly at 00, leading to $E_{0}$ no longer holding the Scaling law. Furthermore, since $v_{eff}(T)$ is only used for non-zero temperatures, that is, $V_{eff}(0)\neq\lim_{T\longrightarrow 0}v_{eff}(T)$.
@@
$\frac{E_{sph}(T)}{T}-7ln\left(\frac{v(T)}{T}\right)+ln\left(\frac{T}{100GeV}\right)>(35.9-42.8).$ (28)
Based on the data tables in Appendix A, the $E/T$ ratio is always greater than $50$; $v_{c}$ is not much bigger than $T_{c}$ and almost $T_{c}$ are larger than $100$ GeV so $[-7\ln(v_{c}/T_{c})+\ln(T/100)]$ has a negative value but not less than $-5$. Thus the sphaleron energy always satisfies the decoupling condition in Eq. (28). Also, the nucleation temperature $T_{N}$ is not much larger than $T_{C}$, the supercooling process is small. In other words, the onset of phase transition is not so slow. However, the smaller $a$ is, the larger the gap between two phases is, the longer the EWPT is, another supercooling process in this model.
$\beta/H^{*}$ is actually quite difficult to calculate, it is estimated by us from the different contributions to the sphaleron energy. Based on numerical calculations, the contribution of gauge field component to this energy is $55\%$ so having an approximation as Eq.(24). In general an approximation, $\frac{\beta}{H^{*}}\approx T\frac{d}{dT}\left(\gamma.\frac{E^{T}_{sph}}{T}\right)$ can be choosed for other models; $\gamma$ depends on the contributions of components.
@@
Our future work follows the results, calculating CP violation, searching and connecting the strength of EWPT, sphalerons or gravitational waves with experimental data.

### v5

@@ I INTRODUCTION
Particle physics today not only studies each individual problem in depth but also combines it with Cosmology and Astronomy. The baryon asymmetry is a problem of intersection among those fields. The possible causes of this problem are summarized and presented quite extensively. Indeed, for Higgs 125 GeV, Standard Model (SM) has only a second order electroweak phase transition when the mass of Higgs is greater than 75 GeV mkn ; SME ; SMEb ; SMEc ; SMEd ; michela . Therefore, SM should be extended (for example, Refs. plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; chr ; cde ; kusenko ). There are 3 main branches: the first is due to the existence of exotic particles (beyond the standard model (SM)); the second is because of parameters or interactions in SM changing at a electroweak transition temperature plv ; 2b ; 2c ; BSM ; BSMb ; majorana ; majoranab ; thdmb ; ESMCO ; elptdm ; elptdma ; elptdmb ; elptdmc ; elptdmd ; phonglongvan ; phonglongvanb ; phonglongvan2 ; SMS ; munusm ; lr ; singlet ; singletb ; singletc ; singletd ; mssm1 ; mssm1b ; mssm1c ; twostep ; twostepb ; twostepc ; 1101.4665 ; 1101.4665b ; 1101.4665c ; jjgb ; jjgc ; jjgd ; scaling1 ; scaling2 ; scaling3 ; Ahriche3b ; Ahriche3c ; Ahriche3d ; Ahriche3e ; chr ; cde ; kusenko ; the third be having more than 4 dimensions of spacetime pn .
@@
  • Furthermore, in Ref. dori ; 124 ; 127 ; kanemura ; thomas there are many scenarios among the additional Higgs bosons were studied in the EWPT problem, specifically $m_{H^{\pm}}=m_{A}$ or $m_{H}=m_{H^{\pm}}$.
@@ II.1 The Higgs potential
$\begin{split}&m_{H^{\pm}}^{2}=-l_{2}v^{2},m_{A}^{2}=-\dfrac{1}{2}(2l_{2}-l_{3})v^{2},\\ &m_{H}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}+\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{H}}{4}.v^{2},\\ &m_{h}^{2}=\dfrac{v^{2}}{4}\bigg{[}2l_{1}+l_{3}-\sqrt{16l_{1}l_{2}(c_{\beta}^{2}-s_{\beta}^{2})^{2}-8l_{2}l_{3}+l_{3}^{2}-4l_{1}l_{3}(c_{\beta}^{4}-6c_{\beta}^{2}s_{\beta}^{2}+s_{\beta}^{4})}\bigg{]}=\frac{f_{h}}{4}.v^{2}\,.\end{split}$ (4)
@@ II.2 The temperature effective Higgs potential and cotronling EWPT
We see that there is a problem in examining the phase transition. Why is $v_{1}$ different from $v_{2}$ (meaning the vacuum breaking process of the two Higgs fields is different) that we examine them together? We examine the electroweak phase transition with two stages, because of the following two reasons: the first, as the above sections pointed out, in 2HDM-$S_{3}$, the masses of particles ($H$ and $h$) can be changed such that there are no mixing terms between $v_{1}$ and $v_{2}$. This approximation is similar to the way of examining a phase transition where $v_{1}$ and $v_{2}$ are both brought to the variable $v$. The second, we do not ignore the mixing of VEVs in the mass components, we only approximate them. Therefore, with $v_{1}$ different from $v_{2}$, this model has two stages of phase transition and assume $v_{1}<v_{2}$, which means $21<a<$ 2 or $1<\tan\beta$. So the approximate effective potential for this model is
$V_{eff}^{S_{3}}\approx V_{eff}^{S_{3}}(v_{1})+V_{eff}^{S_{3}}(v_{2}).$ (5)
The mixing components between $v_{1}$ and $v_{2}$ have been absorbed into $V^{S_{3}}_{eff}(v_{1})$ and $V^{S_{3}}_{eff}(v_{2})$ when using $\tan{\beta}$ or $a$, as shown in Table 2. Hence in Eq.5, although the effective potentials are separated, the mixing components are still present in them. In addition, although the electroweak phase transition is investigated with two phases, these phases may interact with each other through the mixing components in the mass components of the particles. This will be investigated further in this paper.
@@
In addition, the calculation of the effective potential with the daisy loops of the heavy Higgs particles in the temperature effective potential is necessary for completeness of the problem. However, we need to see that the main driving force of these particles is the trigger for the EWPT process. Therefore, adding these loops may complicate the main calculation because their contributions are negligible in the context of the EWPT calculation.
From many conclusions, we propose a way to investigate or control the Higgs potential to have a first order phase transition or $S=\frac{v_{c}}{T_{c}}>1$ ($v_{c}$ and $T_{c}$ are VEV and critical temperature respectively): the first is to build the high temperature effective potential. The second is to select the investigation variable, usually the masses of new particles. Finally, survey in the absence of daisy loops. Specify the range of values for variables to have $S>1$. If the masses of new particles are greater than 200 GeV, there are no need to use additional daisy loops.
@@
According to Fig. 2, the two potential functions corresponding to $v_{2}$ and $v_{1}$ get closer to each other as $a$ increases. With the uniform convergence of second minima (i.e. always less than $246$ GeV), the two effective potentials are always similar to each other. Although Fig.2 only corresponds to one mass pair, it is quite general, other mass pairs are presented in Appendix A.
@@
a $v_{c}$ [GeV] $T_{c}$ [GeV] $S$ $E_{sph,b}^{T}$ [GeV] $\alpha$ $\frac{\beta}{H*}$ $\Omega h^{2}(f_{peak})\times 10^{-14}$
@@ III.2 Gravitational waves
The next, in Table 4, the contribution of the Higgs field component to the total Sphaleron energy is on average about $45\%$ and by using a possible approximation phong2022 , because when calculating GW we only consider the effective potential and scalar field component 71 ; phu5gw ,
$\frac{\beta}{H^{*}}\approx\left[T\frac{d\left(0.45\frac{E^{T}_{sph,b}}{T}\right)}{dT}\right]_{T=T_{C}},$ (24)
@@
a $E^{T}_{Higgs,b}/E^{T}_{sph,b}(\%)$
$(m_{H^{\pm}}(v_{2}),m_{H}(v_{2}))$ $(m_{H^{\pm}}(v_{2}),m_{H}(v_{2}))$ $(m_{H^{\pm}}(v_{2}),m_{H}(v_{2}))$ $(m_{H^{\pm}}(v_{2}),m_{H}(v_{2}))$
$(600,300)[GeV]$ $(500,250)[GeV]$ $(350,160)[GeV]$ $(300,150)[GeV]$
1.1 43.7192 44.1598 45.7986 46.6429
1.2 43.6772 44.0209 45.5623 46.4284
1.3 43.6652 43.9105 45.3493 46.2181
1.4 43.6798 43.8245 45.1576 46.0172
1.5 43.7177 43.7595 44.9851 45.8281
1.6 43.7761 43.7129 44.8297 45.6517
1.7 43.8521 43.6829 44.6897 45.488
1.8 43.9425 43.6677 44.5634 45.3363
1.9 44.0446 43.6661 44.4493 45.1959
2 44.1555 43.677 44.3465 45.0659
Table 4: The ratio between the Higgs component and the total sphaleron energy.
In Table 4, $E^{T}_{Higgs,b}$ is the components containing only the function $h_{b}(\xi)$ in Eq.11, in most cases $E^{T}_{Higgs,b}/E^{T}_{sph,b}$ decreases as $a$ increases. It ranges from $43.67\%$ to $46.64\%$. So the average of this ratio is around $45\%$. We therefore approximate it as Eq.24. We also find that this ratio holds for most models because the contribution of the effective potential component to sphaleron is quite small. Indeed, this ratio in the THD model is exactly the same as the result in Ref. phong2022 .
@@
The gravitational wave detection frequency is solved in the range below a few mHz. At such low frequencies there is currently no experimental data. However, future detectors may be able to detect gravitational energy densities in the small frequency range as summarized in Table 5. The LISA and DECIGO detectors can pick up the gravitational wave signal generated by the EWPT process in the 2HDM-$S_{3}$ model as shown in Figs. 5, 6.
@@
According to Fig. 6, when $f\leq 1$ mHz the gravitational wave results are consistent with DECIDO but when $f\sim$ a few mHz, LISA gives better results.
@@ IV CONCLUSION AND OUTLOOKS
According to the data tables, at the same value of $a$, the ratio $E^{T}_{sph,b}/v_{b}$ is always equal. By the linear scaling law of sphaleron energy scaling1 ; scaling2 ; scaling3 , $E^{T}_{sph,b}/v_{b}=E_{b0}/v_{b0}=const$ ($E_{b0}$ is the sphaleron energy at $T=0$, Eq. (13) calculated by replacing the effective potential at temperature $T$ with one at 0), the evolution (increase or decrease of sphaleron energy) in the two phase transitions is the same. However, $a$ and this progression increase proportionally.
Additionally to test the scaling law (however, this was initially only an approximation to solve the Sphaleron energy problem scaling2 ; scaling3 ). We estimate some cases like Table 6, $E^{T}_{sph,b}/v_{b}$ and $E_{b0}/v_{b0}$ differ by no more than $24\%$. Thus the results are not very compatible with the scaling law or this law is only true in a non-zero temperature range but is not true in the temperature range from 0 to $T_{c}$. For some other cases the calculation also breaks this law as in Ref.phascaling . Because there are many triggers outside SM and energy scales other than SM. Furthermore, the solutions $h(\xi)$ and $f(\xi)$ in Eqs. (13) will approach $1$ very slowly at 00, leading to $E_{0}$ no longer holding the Scaling law. Furthermore, since $v_{eff}(T)$ is only used for non-zero temperatures, that is, $V_{eff}(0)\neq\lim_{T\longrightarrow 0}v_{eff}(T)$.
@@
$\frac{E_{sph,b}^{T}}{T}-7ln\left(\frac{v(T)}{T}\right)+ln\left(\frac{T}{100GeV}\right)>(35.9-42.8).$ (28)
Based on the data tables in Appendix A, the $\frac{E^{T}_{sph,b}}{T}$ ratio is always greater than $50$; $v_{c}$ is not much bigger than $T_{c}$ and almost $T_{c}$ are larger than $100$ GeV so $[-7\ln(v_{c}/T_{c})+\ln(T/100)]$ has a negative value but not less than $-5$. Thus the sphaleron energy always satisfies the decoupling condition in Eq. (28). Also, the nucleation temperature $T_{N}$ is not much larger than $T_{C}$, the supercooling process is small. In other words, the onset of phase transition is not so slow. However, the smaller $a$ is, the larger the gap between two phases is, the longer the EWPT is, another supercooling process in this model.
$\beta/H^{*}$ is actually quite difficult to calculate, it is estimated by us from the different contributions to the sphaleron energy. Based on numerical calculations, the contribution of gauge field component to this energy is $55\%$ so having an approximation as Eq.(24). In general an approximation, $\frac{\beta}{H^{*}}\approx T\frac{d}{dT}\left(\gamma.\frac{E^{T}_{sph,b}}{T}\right)$ can be choosed for other models; $\gamma$ depends on the contributions of components.
@@
Our future work follows the results, calculating CP violation, searching and connecting the strength of EWPT, sphalerons or gravitational waves with experimental data. Furthermore, we will extend these results in combination with some nonstandard cosmology scenarios and dark matter as in Refs.loc1 ; loc2 .